# Equations involving the modular $j$-function and its derivatives

### v1

@@ Abstract.
VA was Supported by Leverhulme Trust Early Career Fellowship ECF-2022-082 (at the University of Leeds) and by EPSRC Fellowship EP/X009823/1 and DKO Fellowship (at the University of Manchester). SE and VM were supported by EPSRC Fellowship EP/T018461/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
The problem of determining which (systems of) equations involving certain classical transcendental functions of a complex variable have solutions is a natural question at the intersection between complex geometry, model theory, and number theory. In complex geometry, it is a form of analytic Nullstellensatz for the given functions; in model theory, it plays an important role in the definability properties of the functions involved; and in number theory, it is related to Schanuel’s conjecture and its analogues (given by special cases of the Grothendieck-André generalised period conjecture). Often the function under consideration is of arithmetic importance. Examples of such classical functions are the exponential functions of semi-abelian varieties and Fuchsian automorphic functions. In this paper we focus on the modular $j$-function and its derivatives.
The first conjecture in this area arose from Zilber’s work on the model theory of complex exponentiation [Zil05, Zil02, Zil15]. It is now referred to as the Exponential (Algebraic) Closedness conjecture or Zilber’s Nullstellensatz. We refer to the general version of the problem as Existential Closedness, or EC for short. An EC conjecture for the $j$-function was proposed in [AK22, §1], and the authors also proposed an extension of the conjecture that includes the derivatives of $j$. This version of EC is often referred to as Existential Closedness with Derivatives, or ECD for short. It states that if the corresponding analogue of Schanuel’s conjecture for $j$ (see [AEK23, §6.3]) does not explicitly predict that a system of equations involving $j$ and its first two derivatives should not have a solution, then it does have a solution. For a precise statement, see [AK22, AEK21, Asl22]. Note that we do not consider the third and higher derivatives of $j$ as these are rational over $j,j^{\prime},j^{\prime\prime}$; see (2.2).
@@
Some special cases of ECD, proving solvability of some simple equations involving just $j^{\prime}$, have been established in [EH21, Gal21]. An ECD statement for “blurrings” (certain multi-valued twists) of $j$ was obtained in [AK22]. All of these papers mostly focused on, and established stronger results for, the $j$-function without derivatives. These results are analogous to their exponential counterparts, namely, [BM17, DFT18, AKM23, Kir19, Zil02, Gal23]. Although different methods have been used across these works, one common feature is that they all exploit in some way the periodicity of $\exp$ or the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-invariance of $j$. Incorporating the derivatives of $j$ into the equations presents then a significant new challenge, as $j^{\prime}$ and $j^{\prime\prime}$ are no longer $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-invariant. It is also worth noting that a differential version of ECD was obtained in [AEK21], and that it is so far the only setting where a full Existential Closedness statement is proved for $j$ with derivatives, and the same method also applies to $\exp$.
In this article we deal with arbitrary equations of one complex variable.
@@
The proof of Theorem 1.1 is based on a generalisation of the method of [EH21] which uses Rouché’s theorem from classical complex analysis to establish some cases of EC for $j$ (without derivatives). This can be seen as an analogue of the classical fact that every irreducible polynomial $p(X,Y)\in\operatorname{\mathbb{C}}[X,Y]$ which depends on $Y$ has infinitely many zeroes of the form $(z,\exp(z))$, unless $p=cY$ for some $c\in\operatorname{\mathbb{C}}^{\times}$.
We can rephrase the problem of solving systems of equations involving $j,j^{\prime},j^{\prime\prime}$ to asking whether an algebraic variety $V\operatorname{\subseteq}\operatorname{\mathbb{C}}^{4n}$ intersects the graph of (Cartesian powers of) the vector-function $(j,j^{\prime},j^{\prime\prime}):\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}^{3}$. For instance, the equation $z+j(z)j^{\prime}(z)+z^{2}j^{\prime\prime}(z)=0$ has a solution if and only if the variety $V\operatorname{\subseteq}\operatorname{\mathbb{C}}^{4}$ defined by $X+Y_{0}Y_{1}+X^{2}Y_{2}=0$ intersects the graph of $(j,j^{\prime},j^{\prime\prime})$. Throughout the paper, all algebraic subvarieties of $\operatorname{\mathbb{C}}^{4}$ will be defined by polynomials in the ring $\operatorname{\mathbb{C}}[X,Y_{0},Y_{1},Y_{2}]$.
@@
Our main goal is to obtain a much stronger version of Theorem 1.1. We show that for any polynomial $F(X,Y_{0},Y_{1},Y_{2})$ the set $\{(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))\in\operatorname{\mathbb{H}}\times\operatorname{\mathbb{C}}^{3}:F(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0\}$ is Zariski dense in the hypersurface $F(X,Y_{0},Y_{1},Y_{2})=0$, unless $F$ is divisible by an explicit (finite) list of polynomials. In this case we say that the equation $F(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$ has a Zariski dense set of solutions (see Definition 2.1), that is, by a solution of such an equation we understand a tuple $(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))$ rather than just $z_{0}$. Proving Zariski density amounts to establishing certain cases of ECD for subvarieties of $\operatorname{\mathbb{C}}^{8}$. For instance, showing that the system $\{j^{\prime\prime}(z)=0,~{}j(z)\neq 0\}$ has a solution is equivalent to showing that the system $\{j^{\prime\prime}(z_{1})=0,~{}j(z_{1})z_{2}=1\}$ has one.
@@
The bulk of the paper is focused on proving the Zariski density of the set of solutions. For instance, the solutions of the equation $zj^{\prime\prime}(z)+(z^{3}+1)j^{\prime}(z)^{2}+j^{\prime}(z)j(z)^{7}=0$ found via the method used to prove Theorem 1.1 are the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-conjugates of $\rho=-\frac{1}{2}+\frac{\sqrt{3}}{2}i$. These are obviously not Zariski dense, for it is well known that $j(\rho)=j^{\prime}(\rho)=j^{\prime\prime}(\rho)=0$. In fact, for Zariski density we need to show at least that the solutions are not contained in finitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits, unless the equation is of the form $j(z)=u$, for some $u\in\operatorname{\mathbb{C}}$.
Observe that
$\forall z\in\operatorname{\mathbb{H}}\left[j(z)(j(z)-1728)=0\Leftrightarrow j^{\prime}(z)=0)\right].$
@@
Thus, the three equations $j(z)=0$, $j(z)-1728=0$, and $j^{\prime}(z)=0$ do not have Zariski dense sets of solutions. Our main result shows that these are essentially the only non-examples.
@@
Theorem 1.2 implies that for every rational function $G(X,Y_{0},Y_{1},Y_{2})\in\operatorname{\mathbb{C}}(X,Y_{0},Y_{1},Y_{2})$, the function $G(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))$ has a zero unless $G$ is of the form
$\frac{Y_{0}^{s}(Y_{0}-1728)^{t}Y_{1}^{l}}{H(X,Y_{0},Y_{1},Y_{2})}$
where $H$ is a polynomial coprime with $Y_{0}(Y_{0}-1728)Y_{1}$, and $s,t,l\in\mathbb{Z}$.
A special case of Theorem 1.2 is that the equation $j^{\prime\prime}(z)=0$ has a Zariski dense set of solutions. [^1] In particular, the ramification points of $j^{\prime}$ are not contained in finitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits. In this case, even proving that there is a solution outside the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbit of $\rho$ is highly non-trivial, see §7.1.
@@
In [Ete22], the author studies the problem of finding generic solutions to equations involving $j$ (and its derivatives). In particular, [Ete22, Theorem 6.5] shows that there is a countable field $C_{j}\subseteq\operatorname{\mathbb{C}}$ such that for any irreducible hypersurface $V\subset\operatorname{\mathbb{C}}^{4}$ satisfying the conditions of Theorem 1.2, if $V$ is not definable over $C_{j}$, then for any finitely generated subfield $K\subset\operatorname{\mathbb{C}}$ over which $V$ can be defined there is a point of the form $(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))\in V$ such that
@@
To prove Theorem 1.2 we establish general criteria for the solvability of certain equations involving periodic functions (see §4). The following is a special case of those criteria.
@@
Let $f_{0},\ldots,f_{n}:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ be $1$-periodic meromorphic functions. Suppose that for some $k$ one of the following conditions is satisfied:
@@
Let us consider an example illustrating how we apply Proposition 1.5 in practice. It also gives an idea of our approach in the general case.
@@
where either $p(j(z))=j(z)(j(z)-1728)$ or $p(j(z))=j(z)^{2}(j(z)-1728)$. First, we want to get an equivalent equation which is written as a sum of powers of $z$ with periodic coefficients. To that end we apply the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-transformation $z\mapsto-\frac{1}{z}$ and, using the identities $j\left(-\frac{1}{z}\right)=j(z),~{}j^{\prime}\left(-\frac{1}{z}\right)=z^{2}j^{\prime}(z)$, we get
@@
When $p(j)=j(j-1728)$, the ratio $\frac{j(j-1728)}{(j^{\prime})^{2}}$ has a pole at $\tau=\rho$. When $p(j)=j^{2}(j-1728)$, the ratio $\frac{j^{2}(j-1728)}{(j^{\prime})^{2}}$ has no finite poles, but it has limit $\infty$ for $\operatorname{Im}(z)\to+\infty$.
Thus, by Proposition 1.5, there is a sequence $z_{m}$ with $z_{m}\to\tau$ and $z_{m}\neq\tau$ in the first case, or $\operatorname{Im}(z_{m})\to+\infty$ and $0\leq\mathrm{Re}(z_{m})\leq 1$ in the second case, such that for all sufficiently large $m$ the point $z_{m}+m$ is a solution to the equation (1.2). This already implies that the solutions of (1.2) intersect infinitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits.
To deduce Zariski density of these solutions, suppose that all of them are also solutions of another independent equation $G(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$. Combining this and (1.2) we can eliminate $z$ and end up with an equation $H(j,j^{\prime},j^{\prime\prime})=0$. Now, our assumption means that $H(j,j^{\prime},j^{\prime\prime})$ vanishes at $z_{m}+m$, hence also at $z_{m}$ by periodicity. This is not possible, for a 1-periodic holomorphic function cannot have infinitely many zeroes with real part bounded from above and below and imaginary part bounded from below. This then implies the Zariski density of solutions of (1.1).
We also note that our criteria can be applied to more general periodic functions, beyond polynomials of $j,j^{\prime},j^{\prime\prime}$, such as $\exp$ or the Weierstrass Weierstrass-$\wp$-function. For instance, Proposition 1.5 implies that the function $j^{\prime}(z)z+\exp(2\pi iz)$ has infinitely many zeroes around the points $i+m$ where $m$ is a large integer.

### v3

@@ Abstract.
VA was Supported by Leverhulme Trust Early Career Fellowship ECF-2022-082 at the University of Leeds (where most of this work was done) and by EPSRC Fellowship EP/X009823/1 and DKO Fellowship at the University of Manchester. SE and VM were supported by EPSRC Fellowship EP/T018461/1 at the University of Leeds. For the purpose of open access, the authors have applied a Creative Commons Attribution ($C$ BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
The problem of determining which (systems of) equations involving certain classical transcendental functions of a complex variable have solutions is a natural question at the intersection between complex geometry, model theory, and number theory. In complex geometry, it is a form of analytic Nullstellensatz for the given functions; in model theory, it plays an important role in the definability properties of the functions involved; and in number theory, it is related to Schanuel’s conjecture and its analogues (given by special cases of the Grothendieck–André generalised period conjecture). Often the function under consideration is of arithmetic importance. Examples of such classical functions are the exponential functions of semi-abelian varieties and Fuchsian automorphic functions. In this paper we focus on the modular $j$-function and its derivatives.
The first conjecture in this area arose from Zilber’s work on the model theory of complex exponentiation [Zil05, Zil02, Zil15]. It is now referred to as the Exponential (Algebraic) Closedness conjecture or Zilber’s Nullstellensatz, and predicts when systems of equations involving addition, multiplication, and complex exponentiation have solutions in the complex numbers. We refer to the general version of the problem as Existential Closedness, or EC for short. An EC conjecture for the $j$-function was proposed in [AK22, §1]; in geometric terms it states that any algebraic variety $V\subseteq\operatorname{\mathbb{C}}^{2n}$ satisfying geometric conditions known as freeness and broadness, intersects the $n$-fold graph of the $j$-function. The definition of these geometric conditions is long and will not be used in the present work, so we refer the interested reader to [AK22, §2.2], but, informally, freeness and broadness ensure that the equations defining $V$ do not break any functional properties of $j$ coming from the linear-fractional action of $\operatorname{GL}_{2}^{+}(\operatorname{\mathbb{Q}})$ (where ++ denotes positive determinant) on the upper half-plane, as well as not contradicting a conjecture on transcendental values of the $j$-function analogous to Schanuel’s conjecture for exponentiation (see [AK22, Conj. 1.1] and [AEK23, §6.3] for the statement of this conjecture).
If one somehow knows that an algebraic variety $V$ does intersect the graph of $j$, a very natural next question is to determine how these intersection points are distributed within $V$. For instance, one may ask whether these points are Zariski dense in $V$. We remark that if $V\subseteq\operatorname{\mathbb{C}}^{2n}$ satisfies the above-mentioned geometric conditions of freeness and broadness, then for any Zariski open subset $V^{\prime}\subseteq V$ it is possible to construct an algebraic variety $W\operatorname{\subseteq}\operatorname{\mathbb{C}}^{2(n+1)}$ which is also free and broad and projects onto $V^{\prime}$. Thus, if we assume EC then, applying it to $W$, we deduce that $V^{\prime}$ intersects the graph of $j$. Since $V^{\prime}$ was an arbitrary Zariski open subset of $V$, we conclude that the intersection of $V$ with the graph of $j$ is Zariski dense in $V$.
In the same work [AK22], the authors also proposed an extension of the conjecture incorporating the derivatives of $j$ (see [AK22, Conj. 1.6]). This version of EC is often referred to as Existential Closedness with Derivatives, or ECD for short. This time the variety $V$ in question is a subset of $\mathbb{C}^{4n}$, and the conjecture states that if $V$ satisfies analogous geometric notions of freeness and broadness (again related to a form of Schanuel’s conjecture, now involving $j$ and its derivatives), then $V$ intersects the $n$-fold graph of the map $z\mapsto(j(z),j^{\prime}(z),j^{\prime\prime}(z))$. For the definitions and precise statements of these conjectures, see [AK22, AEK21, Asl22]. Note that we do not consider the third and higher derivatives of $j$ as these are rational over $j,j^{\prime},j^{\prime\prime}$ (see (2.2)). As with EC, ECD implies that if $V\subseteq\operatorname{\mathbb{C}}^{4n}$ satisfies the geometric conditions of freeness and broadness, then $V$ has a Zariski dense set of points of the desired form.
@@
Very few cases of ECD have been proven, in comparison to EC where various families of varieties in $\operatorname{\mathbb{C}}^{2n}$ have been shown to satisfy the conjecture. Prior to the present work, only very special cases of ECD had been solved, proving solvability of some simple equations involving just $j^{\prime}$ (so not combining it with $j$ or $j^{\prime\prime}$), see [EH21, Gal25]. An ECD statement for “blurrings” (certain multi-valued twists) of $j$ was obtained in [AK22]. All of these papers mostly focused on, and established stronger results for, the EC conjecture for the $j$-function (without derivatives). These results are analogous to their exponential counterparts, namely, [BM17, DFT18, AKM23, Kir19, Zil02, Gal23]. Although different methods have been used across these works, one common feature is that they all exploit in some way the periodicity of $\exp$ or the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-invariance of $j$. Incorporating the derivatives of $j$ into the equations presents then a significant new challenge, as $j^{\prime}$ and $j^{\prime\prime}$ are no longer $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-invariant. It is also worth noting that a differential version of ECD was obtained in [AEK21], and that it is so far the only setting where a full Existential Closedness statement is proved for $j$ with derivatives, and the same method also applies to $\exp$.
In this article we prove the ECD conjecture when $n=1$, which precisely states that any algebraic variety $V\subseteq\mathbb{C}^{4}$ of dimension 3 without constant coordinates contains (a Zariski dense set of) points of the form $(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))$. This amounts to checking exactly which equations of one complex variable involving only $z,j(z),j^{\prime}(z),j^{\prime\prime}(z)$ have solutions, and whether these solutions are Zariski dense. Note that for $n=1$ broadness of $V$ just means $\dim V\geq 3$ hence the only non-trivial case is $\dim V=3$. On the other hand, freeness for $n=1$ means $V$ has no constant coordinates. Nevertheless, we will even be able to decide what happens when $V$ does have a constant coordinate.
Our first main result establishes the existence of solutions in all non-trivial cases.
@@
The proof of Theorem˜1.1 is based on a generalisation of the methods of [EH21], which use Rouché’s theorem from classical complex analysis to establish some cases of EC for $j$ (without derivatives). Theorem˜1.1 can be seen as an analogue of the classical fact that every irreducible polynomial $p(X,Y)\in\operatorname{\mathbb{C}}[X,Y]$ which depends on $Y$ has infinitely many zeroes of the form $(z,\exp(z))$, unless $p=cY$ for some $c\in\operatorname{\mathbb{C}}^{\times}$.
Throughout the paper, all algebraic subvarieties of $\operatorname{\mathbb{C}}^{4}$ will be defined by polynomials in the ring $\operatorname{\mathbb{C}}[X,Y_{0},Y_{1},Y_{2}]$.
Our main goal is to obtain a much stronger version of Theorem˜1.1. We show that for any polynomial $F(X,Y_{0},Y_{1},Y_{2})$ the set
@@
is Zariski dense in the hypersurface $F(X,Y_{0},Y_{1},Y_{2})=0$, unless $F$ is divisible by an explicit (finite) list of polynomials. In this case we say that the equation $F(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$ has a Zariski dense set of solutions (see Definition˜2.1), that is, by a solution of such an equation we understand a tuple $(z_{0},j(z_{0}),j^{\prime}(z_{0}),j^{\prime\prime}(z_{0}))$ rather than just $z_{0}$.
We remind the reader that this is equivalent to establishing certain cases of ECD for subvarieties of $\operatorname{\mathbb{C}}^{8}$: given a hypersurface $V\subseteq\operatorname{\mathbb{C}}^{4}$ and a Zariski open dense $V^{\prime}\subseteq V$, there is $W\subseteq\operatorname{\mathbb{C}}^{8}$ free and broad which projects onto $V^{\prime}$, such that $W$ intersects the graph of $j$ and its derivatives if and only if $V^{\prime}$ does. For instance, the system $\{j^{\prime\prime}(z)=0,~j(z)\neq 0\}$ has a solution if and only if $\{j^{\prime\prime}(z_{1})=0,~j(z_{1})z_{2}=1\}$ does.
@@
The bulk of the paper is focused on proving the Zariski density of the set of solutions, which the proof of Theorem˜1.1 does not provide. For instance, the solutions of the equation $zj^{\prime\prime}(z)+(z^{3}+1)j^{\prime}(z)^{2}+j^{\prime}(z)j(z)^{7}=0$ found via the proof Theorem 1.1 are the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-conjugates of $\rho=-\frac{1}{2}+\frac{\sqrt{3}}{2}i$. These are obviously not Zariski dense, for it is well known that $j(\gamma\rho)=j^{\prime}(\gamma\rho)=j^{\prime\prime}(\gamma\rho)=0$ for every $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ (see [Lan87, p. 40]). Indeed, Zariski density requires at least that the solutions are not contained in finitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits, except for when the equation is of the form $\prod_{k}(j(z)-u_{k})=0$ for some $u_{k}\in\operatorname{\mathbb{C}}$.
The zeroes of $j^{\prime}$ are in fact problematic: observe that
$\forall z\in\operatorname{\mathbb{H}}\left[j(z)(j(z)-1728)=0\Longleftrightarrow j^{\prime}(z)=0\right].$
@@
This immediately gives that the three equations $j(z)=0$, j(z)1728=0j(z)-1728=0, and $j^{\prime}(z)=0$ do not have Zariski dense sets of solutions. Our main result shows that these are essentially the only non-examples.
@@
Theorem˜1.2 implies that for every rational function $G(X,Y_{0},Y_{1},Y_{2})\in\operatorname{\mathbb{C}}(X,Y_{0},Y_{1},Y_{2})$, the function $G(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))$ has a zero unless $G$ is of the form
$\frac{Y_{0}^{s}(Y_{0}-1728)^{t}Y_{1}^{\ell}}{H(X,Y_{0},Y_{1},Y_{2})}$
where $H$ is a polynomial and $s,t,\ell\in\mathbb{N}$.
A special case of Theorem˜1.2 is that the equation $j^{\prime\prime}(z)=0$ has a Zariski dense set of solutions. [^1] In particular, the ramification points of $j^{\prime}$ are not contained in finitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits. In this case, even proving that there is a solution outside the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbit of $\rho$ is highly non-trivial, see §7.1.
@@
In [Ete25], the author studies the problem of finding generic solutions to equations involving $j$ (and its derivatives) under the assumption that the system has a Zariski dense set of solutions. In particular, combining Theorem˜1.2 with [Ete25, Theorem 6.5] we get that there is a countable field $C_{j}\subseteq\operatorname{\mathbb{C}}$ such that for any irreducible hypersurface $V\subset\operatorname{\mathbb{C}}^{4}$ satisfying the conditions of Theorem˜1.2, if $V$ is not definable over $C_{j}$, then for any finitely generated subfield $K\subset\operatorname{\mathbb{C}}$ over which $V$ can be defined there is a point of the form $(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))\in V$ such that
@@
To prove Theorem 1.2 we establish general criteria for the solvability of certain equations involving periodic functions (see §4). The following proposition is a special case of those criteria.
Definition 1.5.
A meromorphic function $f:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ is 1-periodic if $f(z+1)=f(z)$ for every $z\in\operatorname{\mathbb{H}}$. Every such function induces a meromorphic function $\tilde{f}(q)$ on the punctured unit disc by performing the change of variable $q=\exp(2\pi iz)$. We say that $f$ is meromorphic at $i\infty$ if $\tilde{f}$ is meromorphic at 0.
@@
Let $f_{0},\ldots,f_{n}:\operatorname{\mathbb{H}}\to\operatorname{\mathbb{C}}$ be $1$-periodic functions which are meromorphic on $\operatorname{\mathbb{H}}\cup\{i\infty\}$. Suppose that for some $k$ one of the following conditions is satisfied:
@@
Let us consider an example illustrating how we apply Proposition˜1.6 in practice. It also gives an idea of our approach in the general case.
@@
where either p(j(z))=j(z)(j(z)1728)p(j(z))=j(z)(j(z)-1728) or $p(j(z))=j(z)^{2}(j(z)-1728)$. First, we want to get an equivalent equation which is written as a sum of powers of $z$ with periodic coefficients. To that end we apply the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-transformation $z\mapsto-\frac{1}{z}$ and, using the identities $j\left(-\frac{1}{z}\right)=j(z),~j^{\prime}\left(-\frac{1}{z}\right)=z^{2}j^{\prime}(z)$, we get
@@
Thus we obtain an equation in a suitable form for using Proposition 1.6, where $f_{4}=(j^{\prime})^{2}$, $f_{3}=f_{2}=f_{1}=0$ and $f_{0}=p(j)$. When p(j)=j(j1728)p(j)=j(j-1728), the ratio $\frac{f_{0}}{f_{4}}=\frac{j(j-1728)}{(j^{\prime})^{2}}$ has a pole at $\tau=\rho$. When $p(j)=j^{2}(j-1728)$, the ratio $\frac{f_{0}}{f_{4}}=\frac{j^{2}(j-1728)}{(j^{\prime})^{2}}$ has no finite poles, but it has limit $\infty$ as $\operatorname{Im}(z)\to+\infty$.
Thus, by Proposition˜1.6, there is a sequence $z_{m}$ with $z_{m}\to\tau$ and $z_{m}\neq\tau$ in the first case, or $\operatorname{Im}(z_{m})\to+\infty$ and $0\leq\mathrm{Re}(z_{m})\leq 2$ in the second case, such that for all sufficiently large $m$ the point $z_{m}+m$ is a solution to the equation (1.2). This already implies that the solutions of (1.2) intersect infinitely many $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbits.
To deduce Zariski density of these solutions, suppose that all of them are also solutions of another independent equation $G(z,j(z),j^{\prime}(z),j^{\prime\prime}(z))=0$. Combining this and (1.2) we can eliminate $z$ and end up with an equation $H(j,j^{\prime},j^{\prime\prime})=0$. Now, our assumption means that $H(j,j^{\prime},j^{\prime\prime})$ vanishes at $z_{m}+m$, hence also at $z_{m}$ by periodicity. This is not possible, for a 1-periodic holomorphic function, meromorphic at $i\infty$, cannot have infinitely many zeroes with real part bounded from above and below and imaginary part bounded from below. This then implies the Zariski density of solutions of (1.1).
We also note that our criteria can be applied to more general periodic functions, beyond polynomials of $j,j^{\prime},j^{\prime\prime}$, such as $\exp$ or the Weierstrass $\wp$-function. For instance, Proposition˜1.6 implies that the function $j^{\prime}(z)z+\exp(2\pi iz)$ has infinitely many zeroes around the points $i+m$ where $m$ is a large integer.